Tech Staff Meeting Notes September 29, 2015

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


Priorities


News & Highlights

dori (the planned folkworm replacement) is now running heli and getting data from nori waveservers.

First aid kits in all three seisnet vehicles were replaced with new OSHA approved kits.


NMSZ

Put this on hold for now. Jim requested the extra sim card now so that it could be used for several noise studies on deck. The six stations in the AO network are being decommissioned and the hardware returned to the ANSS Depot. CERI offered to take over O&M for two: U40A (near Mountain Home, AR) and X40A (east of Hot Springs). Jim did a recon on U40A and its doable but will take a little work. This effort will likely take about 6 months.

2015 Preventive Maintenance78% complete
NM BB/SM & SP: 63 done & 14 to go
NM SM: 9 done & 8 to go
AG BB/SM: 5 done & 1 to go
GPS: 13 done & 0 to go
Misc: 4 done & 4 to go

GSAR is back online at Gosnell and we're waiting for the new Cisco firewall device scheduled to be installed on September 21st to revisit the UDP packet problem. This will put the school on the Arkansas State Network, similar to the AGS stations, so hopefully we'll have better success with the UDP packet problem.

freewave serial reset installed at 9 of 11 broadbands.

We have 11 episensors cannabalized from the K2's. We'll replace the RT147's at all 11 NM broadband stations then cull the RT147's for working sensors. So far PEBM, GNAR, HBAR, PENM, LNXT, GLAT, HENM, HALT, and HICK are complete. LPAR and PARM remain.

NMDown

Down
HOVM.EHN see image

Misc Problems
CPAR/TYAR - spikes
LPAR.HNN drifting see image

Two ISIS3 upgrades remain: ARPT, and the PLAR replacement.

RDST has a very large pole right next to the vault that will corrupt any data during strong ground motion. It needs to get moved; Dave has a couple options on the Rhodes campus (the name will change when its moved). Hopefully rather than reinstall the Guralps in the new location, we can install a Basalt instead.

NODE Tasks

The win7 upgrades at nodes (mudworm is done).

DBB: 900 MHz link between Wilder & SFTN failed. We're working on guying the tower on Wilder Towers and should have a 4' grid parabolic dish installed by weeks end.

James reported that the DBB link between CERI and Richardson periodically goes down for about a minute. It comes back automatically and everything seems to catch up ok, but real-time packets are lost. This issue may be resolved by switching to Wilder; Nope, the bottleneck is at SFTN.

MKTA:

LNXT:

NMAD:

The waveserver continues to have occasional very gappy periods most likely due to missed messages. Steve will investigate. Example is here.

There were daily gaps in the data of 2 to 3 hours at about 2 or 3 AM for the NMAD broadband stations. James correct the problem by resettin the nmad and lnxt aironets. Are one or both of these heading for failure?

AZTN/HDBR: Shahram submitted a proposal to TDOT to replace all the sensors with Episensors (including new cabinets, etc). We asked USGS to provide new radios so we can have a continuous buffer of all channels at AZTN. We'll then allow access to NSMP so they can extract event data. Unofficial word is TDOT approved the funding.

CERI:

Need to review and update the power systems in 218.


ETSZ

Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Need to do a noise study.

ETDown

Down
MTTN Node- ? (propable issue with NetTN circuit down).

Misc. Problems
CPRT.HNN see image
VHTN.EHE see image
VHTN.EHN see image
ASTN.HNx

STAR Node

Improve RCGA AC power and perform maintenance work on shed.

Replace BCRT Strong Motion Vault and instrument. Improve transient suppression/grounding.

Replace GMG RT-147.

Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are too tall to get a signal out. Looking for an alternate site.

BRBC Node

Morgantown node hardware is ready to go (needs an earthworm) and waiting on completion of the ice bridge and fiber. Delivery of ice bridget and cable rack is June 23 so we're probably looking at an October node install. Still waiting on completion of the facility.

Upgrade all High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, SMNC, TVNC, TRYN.

GFM needs to be upgraded to ISIS4. If we can get a helicopter drop for equipment and supplies then we'll also upgrade to 3-c S-13. Should also get the helicopter to remove old hardware at the same time.

MTTN Node

Replace VHTN E-W and N-S Instruments.

WMTN pit needs to be moved due to erosion.

CCNC is not going through fgtn and instead repeats through ASTN. The 950MHz shot would work so it could be converted to digital.

We're still repeating CPRT through GTTN. The new internet host at the Corryton Senior Center for CPRT has been verbally approved by Greg Hill the center coordinator and the Knox County facility supervisor.


GPS

CJTR comm is down.


AG Network

HHAR needs a new router and power supply.

WHAR replace RT147 with episensor

FCAR AC power upgrade.


FDL

The FDL wiki is here: https://umwiki.memphis.edu/display/FDL/CERI+FDL


MISCELLANEOUS

The seisnet group on umdrive is at http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password.